Use the following information for Chambers Corporation to answer the following questions:

                          20x1                                        20x2

Net Income         $825,000                                $930,000

Depreciation Expense        130,000                                    165,000

Inventory                             1,200,000                               1,640,000

Gross Fixed Assets             4,500,000                               4,850,000

Dividends Paid                    110,000                                   110,000

Accounts Receivable           700,000                                   810,000

Issued Common Stock        1,500,000                                0

Accounts Payable                400,000                                430,000

Long-Term Debt                   1,100,000                            1,600,000

Accruals                               150,000                               140,000

Net Fixed Assets                  2,000,000                            2,550,000

Cash                                    400,000                                1,005,000

Calculate the net cash from operating activities for Chambers Corporation for the year 20x2?

Select one:

a.

​Negative $260,000

b.

​$1,170,000

c.

​$565,000

d.

​$85,000
Answer #1

cash flow from operating activities = net income + depreciation - increase in inventory - increase in account receivables + increase in account payables - decrease in accruals

  = 930000+165000-(1640000-1200000)-(810000-700000)+(430000-400000)-(150000-140000)

= 930000+165000-440000-110000+30000-10000 = 565000